given the function f(x)= 3 abs(x-6) calculate the following values
1.) f(x+1)
2.) f(x^2+2)

Answer by Theo(2973)

f(x) = 3*abs(x-6)
-----
f(x+1) = 3*abs((x+1)-6)
This becomes:
f(x+1) = 3*abs(x-5)
-----
f(x^2+2) = 3*abs((x^2+2)-6)
This becomes:
f(x^2+2) = 3*abs(x^2-4)
-----
in the first case, you replaced x with (x+1) in the equation.
in the second case, you replaced x with (x^2+2) in the equation.
-----